# Assessing survival models by interval testing

### v1

@@ Abstract
When considering many survival models, decisions become more challenging in health economic evaluation. In this paper, we present a set of methods to assist with selecting the most appropriate survival models. The methods highlight areas of particularly poor fit. Furthermore, plots and overall p-values provide guidance on whether a survival model should be rejected or not.
Keywords: Survival analysis, calibration, health economics, decision making
@@ 1 Introduction
The primary results of a clinical trial are often right-censored survival data. To understand the efficacy of a treatment, survival models must be fitted to this data. A choice between fitted survival models must then be made, and this is often the primary driving factor in the economic evaluation of a new medical product. During this process, a large number of models are usually considered. However, with a growing emphasis on flexible survival mdoels[1] and a requirement for standard parametric models[2], the number of models considered can easily exceed 20, making decisions challenging. Our proposed methods can assist this process by indicating which survival models provide implausible estimates.
@@
When we plot these survival models, we will refer to them as survival curves. Recommendations on selecting the most appropriate survival models are provided in NICE TSD 14[2]. These include visual curve inspection, AIC/BIC tests, clinical validity, and the use of external data. Except for AIC/BIC, these approaches are subjective and often time-consuming to implement, especially when considering many curves. It would be useful if curves were quickly shown to be implausible. This would reduce the number of feasible curves, making the decision process easier. One approach that may facilitate this is calibration, but this is rarely used in health technology assessment.
@@
In survival analysis, calibration refers to the agreement between predicted probabilities and observed events within a specific time interval. The most common calibration methods are extensions of the Hosmer-Lemeshow goodness-of-fit test[3],[4]. These include the Grønnesby and Borgan test (specific to the Cox proportional hazard model)[5], the Nam and D’Agostino test (applicable to more general models)[6], as well as smoothed calibration curves and their associated metrics[7]. In all of these approaches, patients are assigned a risk score based on the fitted model and are grouped (often into deciles) based on these scores. The groups are then evaluated to determine how well-calibrated a survival model is. Our approach is different from this because patients do not need to be assigned risk scores. This is especially useful for models where the risk scores are equal among all patients, which is the case for our worked examples.
@@
The paper is structured as follows: methods, application, and discussion. The methods section describes: a new calibration approach with intervals based on censoring, an extension to this where intervals can be specified, and finally methods for generating an overall p-value. The application section considers two examples. The first example evaluates a simple exponential model, providing an introduction to the developed methods. The second example is more practical and considers survival data which is challenging to model. In this example, seven parametric survival models are fitted. Subsequently, the full range of developed methods evaluate which curves are most appropriate. Finally, before the discussion is given, the results of a simulation experiment are summarized.
@@ 2.1 Interval testing
We consider right-censored survival data, where for each subject we have time and event given by $(t,E)\in(\mathbb{R}_{>0},\{1,0\})$. We split this data into two types:
@@
Subsequently, we define the set of unique censor times $t_{j,0}^{\prime}$ for $j=1,\ldots,J$. That is, the values $t_{j,0}^{\prime}$ are the unique values from the set of $t_{i,0}$ values, and thereby the $t_{j,0}^{\prime}$ values satisfy $0<t_{j,0}^{\prime}<t_{j+1,0}^{\prime}$. We now define the intervals of sequential censor times such that for $i=1,...,J-1$
$\begin{split}I_{0}&=(0,\ t_{1,0}^{\prime}]\\ I_{i}&=(t_{i,0}^{\prime},\ t_{i+1,0}^{\prime}]\\ I_{J}&=(t_{J,0}^{\prime},\ \infty).\end{split}$
We are interested in the hypothesis that the event generating process is described by the model $M(\boldsymbol{\beta})$, with probability density function $f_{M}$ and fixed parameters $\boldsymbol{\beta}$.
Thereby, imagine we have not seen the data. However, we still have the intervals $I_{j}$ and the model $M(\boldsymbol{\beta})$. We now make the assumption that the intervals themselves, $I_{j}$, are uninformative of the event generating process. It follows that each patient that enters the interval $I_{j}$:
  • either experiences an event or does not experience an event in interval $I_{j}$.
  • has the same probability of experiencing an event as any other patient entering $I_{j}$.
  • has a probability of experiencing an event defined by the event generating process $M(\boldsymbol{\beta})$ (when assuming the null hypothesis holds).
It follows that under the null, the following process occurs within each interval:
@@
where $events(I_{j})$ are the number of events observed within the interval $I_{j}$.
@@
[…]
Finally, $N_{I_{j}}=$“the number of subjects at the start of interval $I_{j}$.” That is, the number of starting patients $(C+E)$, minus the patients lost to events in the previous $j-1$ intervals, minus the patients lost to censors in the previous $j-1$ intervals. That is,
$N_{I_{j}}=\ C+E\ -\sum_{k=0}^{j-1}events(I_{k})\ -\sum_{k=0}^{j-1}censors(I_{k}),$ (3)
where we define the number of censors observed within the interval $I_{j}$ as $censors(I_{j})$. Note that this quantity usually takes a value of 1. (However, there are exceptions to this. For example, the final interval $J$ has $censors(I_{J})=0$ and intervals with multiple censors of equal time at the upper boundary of the interval have $censors(I_{j})>1$.)
For each interval $I_{j}$, we have specified a binomial model with known $N_{I_{j}}$, $p_{I_{j}}$, and observed $events(I_{j})$. Each interval $I_{j}$ thereby has a midpoint p-value for observing $events(I_{j})$ in that interval.
Extension to specified intervals
The methodology above describes how events can be modelled within each interval. However, the intervals themselves are entirely determined by the censor times. It is often of interest to understand events in specified intervals (for example 10 evenly-spaced intervals up to the last censor.) Below, we extend the methodology to allow for such an analysis.
@@
$\begin{split}V_{1}&=(S_{0},\ S_{1}]\\ ...\\ V_{k}&=(S_{k-1},\ S_{k}].\end{split}$
We have unique censor times $t_{j,0}^{\prime}$ for $j=1,\ldots,J$. Define the set of unique times $\tau_{i}$ as the unique set of the union of the $t_{j,0}^{\prime}$ values and the $S_{k}$ values, with $\tau_{i}<\tau_{i+1}$ for $i=1,...,L_{\tau}-1$, where $L_{\tau}$ denotes the number of unique $\tau$ values. We now construct the intervals
$\begin{split}I_{0}^{\prime}&=(0,\tau_{1}]\\ ...\\ I_{i}^{\prime}&=(\tau_{i},\tau_{i+1}]\\ ...\\ I_{L_{\tau}}^{\prime}&=(\tau_{L_{\tau}},\infty].\end{split}$
@@
with $p_{I_{i}^{\prime}}$ and $N_{I_{i}^{\prime}}$ defined analogously to Equations 2 and 3, with the sum in Equation 3 now being over the $I^{\prime}$ intervals.
We now consider the $I_{i}^{\prime}$ intervals as the subintervals that make up each $V_{k}$ interval. Define the collection of subintervals $l=1,...,L_{k}$ as
$\{I_{l,V_{k}}^{\prime}\}=\{I_{i}^{\prime}\ \ \ such\ that\ \ I_{i}^{\prime}\subseteq V_{k}\}.$
@@
$\begin{split}events(I_{1,V_{k}}^{\prime})&\sim binomial(N_{I_{1,V_{k}}^{\prime}},p_{I_{1,V_{k}}^{\prime}})\\ ...\\ events(I_{L_{k},V_{k}}^{\prime})&\sim binomial(N_{I_{L_{k},V_{k}}^{\prime}},p_{I_{L_{k},V_{k}}^{\prime}}).\\ \end{split}$
Thereby $\sum^{L_{k}}_{l=1}events({I_{l,V_{k}}^{\prime}})$ is distributed as the realization of a sum of binomials with known parameters. That is, the number of observed events in each interval $V_{k}$ follows a known distribution. We can thus obtain a midpoint p-value for each specified interval $V_{k}$.
@@ 2.2 Test statistics
The primary recommendation of this paper is the use of interval plots, such as Figure 1b. To support these plots, a single overall p-value can be a useful tool. Such a p-value can be obtained by condensing the plot’s constituent p-values into a single overall p-value.
When deriving these test statistics, we assume that the midpoint p-value for each interval follows a continuous uniform[0,1] distribution under H0. When it comes to the examples, this assumption does not necessarily hold. This is justified in detail in Appendix A and C.
@@
2.2.1 Transformed Fisher test (TFT) statistic
@@
For each interval, we have a midpoint p-value, $p_{j}$. Under H0, we make the assumption these p-values are $uniform[0,1]$. The p-values are transformed to the form
@@
2.2.2 Protection Against Very Small Intervals (PAVSI) test statistic
@@
For each interval, we have a midpoint p-value, $p_{j}$. Under H0, we make the assumption these p-values are $uniform[0,1]$. The p-values are transformed to the form
$V_{j}=\begin{cases}1\ \ \ for\ \ p_{j}\geq 0.975\ \ or\ \ p_{j}\leq 0.025\\ 0\ \ \ else\end{cases}$
It follows that under H0 we have $V_{j}\sim bernoulli(0.05)$, and thereby
$T_{pavsi}=\sum_{j=1}^{I}V_{j}\sim binom(I,0.05),$
where we reject for large $T_{pavsi}$ values. It is noted that $T_{pavsi}$ is a discrete test statistic, and we recommend reporting the midpoint p-value for this, see Appendix A.
Compared to the transformed Fisher test, the PAVSI is more pragmatic in terms of extreme p-values. However, for PAVSI, a p-value of say 0.026 would contribute equivalently to a p-value of 0.5, which is arguably a limitation of this approach.
@@ 3 Application
In example 1, we focus on the first approach only. In the second example, both approaches are described. Note that the two test statistics (2.2) can be used in either approach.
For our examples, we use real world overall survival data for melanoma as is described in [10]. Example 1, uses the ‘Dabrafenib’ treatment arm of the BREAK-3 trial[11]. Example 2, uses the ‘Dabrafenib+Trametinib’ treatment arm of the COMBI-d trial[12]. All data and code is available on Github here.
@@ 3.1 Example 1: A fitted exponential model
To perform the test, we obtain the values $I_{j}$, $N_{I_{j}}$, and $events(I_{j})$ directly from the data. From Equation 1, it remains to obtain $p_{I_{j}}$ for the fitted exponential model. As the probability density function for the exponential distribution is given by $f_{M}(t)=\lambda e^{-\lambda t}$, using Equation 2, it follows that
@@
As the value $\lambda$ is known (from fitting the exponential model to the data), it follows that we have fully specified binomial realisations for each interval $I_{j}$. This information for the first five intervals is presented in Table 1.
$I_{j}$ N.risk $p_{I_{j}}$ $events(I_{j})$ $\mathbb{E}$(events) p-mid p-rand
(0, 1] 187 0.026 1 4.8 0.027 0.034
(1, 1.6] 185 0.016 1 3 0.127 0.111
(1.6, 2.5] 183 0.024 1 4.4 0.037 0.059
(2.5, 3.4] 181 0.023 4 4.2 0.491 0.538
(3.4, 4.2] 176 0.021 5 3.6 0.771 0.762
Table 1: Summary of interval data for Example 1
@@
The p-value for each interval is discrete as it comes from a binomial model. It follows that each discrete p-value does not follow a continuous uniform[0,1] distribution under H0. This can be rectified by using randomized p-values[13]. However, in the interest of reproducibility and clarity, we will focus our discussion on the results of the midpoint p-values. (This is discussed in more detail in Appendix A.)
Figure 1: Part a: The KM curve for the data in purple, censors are given by dashed vertical black lines, and the fitted exponential curve is the overlaid curve in black. Part b: the binomial interval test for this data. Intervals are defined by the censors, where for each interval, red indicates a Bonferroni rejection, and grey indicates an individual flag. Part c: Numbers at risk for the data
@@
For each interval, we consider the midpoint p-value with respect to a 2-sided test. That is, if $p_{I_{j}}\leq 0.025$ or $p_{I_{j}}\geq 0.975$, we have reason to suspect the interval generating said value. We call this the individual test.
@@
Of greater concern, are the intervals that fail the Bonferroni test[14]. The Bonferroni test is chosen to strongly control the familywise error rate at a level of 0.05. In particular, we reject if $p_{I_{j}}\leq 0.025/I$ or $p_{I_{j}}\geq 0.975/I$, where $I=$“the total number of intervals”. This means that under H0 (that the model holds for all intervals) a single Bonferroni rejection of an interval is only expected to occur for 1 in 20 times the data is generated under H0.
@@
In Figure 1, we present the results for the Bonferroni and individual tests alongside the Kaplan-Meier (KM) curve, the fitted exponential model, and the numbers at risk. A single Bonferroni rejection of an interval is good reasoning for rejecting a model. In Figure 1, the extremely poor fit (Bonferroni rejection) is at the end of the curve. This is very unfavourable in health economic evaluation, where plausible extrapolations beyond the observed data are paramount to model selection. That is, the Bonferroni rejection at the end of the curve indicates that the extrapolation will be poor, which is more detrimental compared with a Bonferroni rejection at the start of the curve.
@@
An overall p-value for the fitted exponential model is also of interest. For the transformed Fisher test statistic, there are $I=43$ intervals, and our test statistic was computed as $-2\sum_{j=1}^{I}log(U_{j})=81.84$, giving a p-value of 0.607. For the PAVSI test statistic, there are $I=43$ intervals. From Figure 1, we can read off that $T_{pavsi}=4$ (as there are 3 individual rejections and 1 Bonferroni rejection). This gives us an overall midpoint p-value of 0.114. (These test statistics should be used with caution. In Appendix C, the test statistics are shown to be highly conservative when using the censor interval approach.)

### v2

@@ Abstract
When considering many survival models, decisions become more challenging in health economic evaluation. In this paper, we present a set of methods to assist with selecting the most appropriate parametric survival models. The methods highlight areas of particularly poor fit. Furthermore, plots and overall p-values provide guidance on whether a parametric survival model should be rejected or not.
Keywords: Survival analysis, model selection, calibration, health economics, decision making
@@ 1 Introduction
Right-censored survival data are frequently collected in clinical trials and often form the basis of a trial’s primary conclusions. To model the efficacy of the treatment within and beyond the duration of such trials, survival models can be fitted to such data. A choice between fitted survival models must then be made, and this is often the primary driving factor in the economic evaluation of a new medical product. During this process, a large number of models are usually considered. However, with a growing emphasis on flexible survival models [1] and a requirement for standard parametric models [2], the number of models considered can easily exceed 20, making decisions challenging. Our proposed methods can assist this process by indicating which survival models provide implausible estimates.
@@
Recommendations on selecting the most appropriate survival models are provided in NICE TSD 14 [2]. These include visual curve inspection, AIC and BIC tests, clinical validity, and the use of external data. Except for AIC and BIC, these approaches are subjective and often time-consuming to implement, especially when considering many curves. It would be useful if certain models were quickly shown to be implausible. This would reduce the number of feasible models, making the decision process easier. One approach that may facilitate this is calibration, but this is rarely used in health technology assessment.
@@
In survival analysis, calibration refers to the agreement between predicted probabilities and observed events within a specific time interval. The most common calibration methods are extensions of the Hosmer-Lemeshow goodness-of-fit test [3], [4]. These include the Grønnesby and Borgan test (specific to the Cox proportional hazard model) [5], the Nam and D’Agostino test (applicable to more general models) [6], as well as smoothed calibration curves and their associated metrics [7]. In all of these approaches, patients are assigned a risk score based on the fitted model and grouped (often into deciles) according to these scores. The groups are then evaluated to determine how well-calibrated a survival model is. Our approach is different from this because patients do not need to be assigned risk scores. This is especially useful for models where the risk scores are equal among all patients, such as in the standard parametric models recommended in NICE TSD 14 [2].
@@
The paper is structured as follows: methods, application, simulation, and discussion. The methods section describes: a new calibration approach with intervals based on censoring, an extension to this where intervals can be specified, and methods for generating an overall p-value. The application section considers two examples. The first example evaluates a simple exponential model, providing an introduction to the developed methods. The second example is more practical and considers survival data which is challenging to model. In this example, seven parametric survival models are fitted, where the full range of developed methods are applied to evaluate which models are most appropriate. Subsequently, a simulation elucidates the validity of the various approaches, and type I error rates are presented for various data scenarios. Finally, a discussion is given.
@@ 2.1 Interval testing
We consider right-censored survival data, where for each subject we have (time, event) $\in(\mathbb{R}_{>0},\{1,0\})$. We split this data into two types:
@@
Subsequently, we define the set of unique censor times $t_{j,0}^{\prime}$ for $j=1,\ldots,J$. That is, the values $t_{j,0}^{\prime}$ are the unique values from the set of $t_{i,0}$ values, and thereby the $t_{j,0}^{\prime}$ values satisfy $0<t_{j,0}^{\prime}<t_{j+1,0}^{\prime}$. We now define the intervals of sequential censor times such that for j=1,2,,Jj=1,2,...,J
$I_{j}=\begin{cases}(0,\ t_{1,0}^{\prime}]\ \ \ \ \ \ \ \ \ for\ \ \ j=1\\ (t_{j-1,0}^{\prime},\ t_{j,0}^{\prime}]\ \ \ \ else\end{cases}$
We are interested in the hypothesis that the event generating process is described by the model $M(\boldsymbol{\beta})$, with probability density function $f_{M}$ and fixed parameters $\boldsymbol{\beta}$. We now make the assumption that the intervals themselves, $I_{j}$, are uninformative of the event generating process. It follows that under the null, the following process occurs within each interval
@@
for cases where $N_{I_{j}}\geq 1$.
$events(I_{j})$ are the number of events observed within the interval $I_{j}$.
@@
$\begin{split}p_{I_{j}}&=\mathbb{P}(T\in I_{j}\ |\ T>min(I_{j})\ )\\ &=\mathbb{P}(T\in I_{j})\ /\ \mathbb{P}(T>min(I_{j}))\\ &=\int_{I_{j}}f_{M}(t)dt\Big/\int_{min(I_{j})}^{\infty}f_{M}(t)dt.\end{split}$ (2)
Finally, $N_{I_{j}}=$“the number of subjects eligible to experience an event at the start of interval $I_{j}$,” which we define for j=2,,Jj=2,...,J as
$\begin{split}&N_{I_{1}}=\ C+E\ -censors(I_{1}),\\ &N_{I_{j}}=\ C+E\ -\sum_{k=1}^{j-1}events(I_{k})\ -\sum_{k=1}^{j}censors(I_{k}),\end{split}$ (3)
where we define the number of censors observed within the interval $I_{k}$ as $censors(I_{k})$.
In summary, for intervals $I_{j}$ we have specified a binomial model with known $N_{I_{j}}$, $p_{I_{j}}$, and observed $events(I_{j})$. Each interval $I_{j}$, where $N_{I_{j}}\geq 1$, thereby has a midpoint p-value for observing $events(I_{j})$ in that interval.
2.2 Extension to specified intervals
The methodology above describes how events can be modelled within each interval. However, the intervals themselves are entirely determined by the censor times. It is often of interest to understand events in specified intervals (for example 10 evenly spaced intervals up to the last censor). Below, we extend the methodology to allow for such an analysis.
@@ 2.2 Extension to specified intervals
$V_{k}=\begin{cases}(S_{0},\ S_{1}]\ \ \ \ \ \ \ \ \ for\ \ \ k=1\\ (S_{k-1},\ S_{k}]\ \ \ \ \ \ else\end{cases}$
where $0<S_{0}<S_{1}<\dots<S_{K}$.
We wish to define a unique set of times $\tau_{i}$. We have unique censor times $t_{j,0}^{\prime}$ for $j=1,\ldots,J$ and unique $S_{k}$ values for $k=0,\dots,K$, taking the unique values of the union of these sets, we obtain the values $0<\tau_{i}<\tau_{i+1}$. From these, we construct the $I^{\prime}$ intervals for $i=1,2,...,L_{\tau}$, where $L_{\tau}$ denotes the total number of $\tau$ values.
$I^{\prime}_{i}=\begin{cases}(0,\tau_{1}]\ \ \ \ \ \ \ \ \ for\ \ \ i=1\\ (\tau_{i-1},\tau_{i}]\ \ \ \ \ \ else\end{cases}$
@@
with $p_{I_{i}^{\prime}}$ and $N_{I_{i}^{\prime}}$ defined analogously to Equations 2 and 3, with all $I$ intervals in these formulae being replaced by $I^{\prime}$ intervals.
We now consider the $I_{i}^{\prime}$ intervals as subintervals that make up each $V_{k}$ interval. For each $V_{k}$, $k=1,\dots,K$, we define the collection of $L_{k}$ subintervals with $m=1,...,L_{k}$ as
$\{I_{m,V_{k}}^{\prime}\}=\{I_{i}^{\prime}\ \ \ such\ that\ \ I_{i}^{\prime}\subseteq V_{k}\}.$
@@
[…]
Thereby $\sum^{L_{k}}_{m=1}events({I_{m,V_{k}}^{\prime}})$ is distributed as the realization of a sum of binomials with known parameters. That is, the number of observed events in each interval $V_{k}$ follows a known distribution. (In our code, we utilise that this is a special case of the Poisson binomial distribution.) We can thus obtain a midpoint p-value for each specified interval $V_{k}$.
@@ 2.3 Test statistics
The primary recommendation of this paper is the use of interval plots, such as Figure 1b. To support these plots, a single overall p-value can be a useful tool. Such a p-value can be obtained by condensing the plot’s constituent p-values into a single overall p-value. When deriving these test statistics, we assume that the p-value for each interval follows a continuous uniform[0,1] distribution under H0.
@@
Transformed Fisher test (TFT) statistic
@@
For each interval, we have a p-value, $p_{j}$. Under H0, we make the assumption these p-values are $uniform[0,1]$. The p-values are transformed to the form
@@
Protection Against Very Small Intervals (PAVSI) test statistic
@@
For each interval, we have a p-value, $p_{j}$. Under H0, we make the assumption these p-values are $uniform[0,1]$. The p-values are transformed to the form
$W_{j}=\begin{cases}1\ \ \ for\ \ p_{j}\geq 0.975\ \ or\ \ p_{j}\leq 0.025\\ 0\ \ \ else\end{cases}$
It follows that under H0 we have $W_{j}\sim bernoulli(0.05)$, and thereby
$T_{pavsi}=\sum_{j=1}^{I}W_{j}\sim binom(I,0.05),$
where we reject for large $T_{pavsi}$ values. It is noted that $T_{pavsi}$ is a discrete test statistic. Therefore, in Examples 1 and 2, the PAVSI is calculated using the midpoint p-value approach.
If the reader is interested in using PAVSI in there own work, there are some technical points to consider that are discussed in Appendix B.
@@ 3 Application
In example 1, we focus on the first approach only. In example 2, both approaches are described. For examples 1 and 2, we use real world overall survival data for melanoma from Freeman et al [10]. Example 1 uses the ‘Dabrafenib’ treatment arm of the BREAK-3 trial [11]. Example 2 uses the ‘Dabrafenib+Trametinib’ treatment arm of the COMBI-d trial [12]. All data and code is available on Github here.
@@ 3.1 Example 1: A fitted exponential model
To perform the test, we first obtain the values $I_{j}$, $N_{I_{j}}$, and $events(I_{j})$ directly from the data. To complete equation 1, we must derive $p_{I_{j}}$ for the fitted exponential model. As the probability density function for the exponential distribution is given by $f_{M}(t)=\lambda e^{-\lambda t}$, using Equation 2, it follows that
@@
The value $\lambda$ is known (from fitting the exponential model to the data). We have therefore fully specified the binomial realisations for each interval $I_{j}$. This information for the first five intervals is presented in Table 1.
$I_{j}$ N.risk $p_{I_{j}}$ $events(I_{j})$ $\mathbb{E}$(events) p-mid
(0, 1] 187 0.026 1 4.8 0.028
(1, 1.6] 185 0.016 1 2.9 0.129
(1.6, 2.5] 183 0.024 1 4.4 0.038
(2.5, 3.4] 181 0.023 4 4.2 0.496
(3.4, 4.2] 176 0.021 5 3.6 0.775
Table 1: Example 1: Summary of interval data.
@@
The p-value for each interval is discrete as it comes from a binomial model. It follows that each discrete p-value does not follow a continuous uniform[0,1] distribution under H0. This can be rectified by using randomised p-values [13]. However, in the interest of reproducibility and clarity, we will focus our discussion on the results of the midpoint p-values. (This is discussed in more detail in Appendix A.)
@@
For each interval, we consider the midpoint p-value with respect to a 2-sided test. That is, if $p_{I_{j}}\leq 0.025$ or $p_{I_{j}}\geq 0.975$, we have reason to suspect the interval generating said value. We call this the individual test. Note that this is an informal tool, and we do accept or reject a model on the basis of a single individual test.
@@
Of greater concern, are the intervals that fail the Bonferroni test [14]. The Bonferroni test is chosen to strongly control the familywise error rate at a level of 0.05. In particular, we reject if $p_{I_{j}}\leq 0.025/I$ or $p_{I_{j}}\geq 0.975/I$, where I=I=“the total number of intervals”. This means that under H0 (that the model holds for all intervals) a single Bonferroni rejection of an interval is only expected to occur for 1 in 20 times the data is generated under H0.
@@
In Figure 1, we present the results for the Bonferroni and individual tests alongside the Kaplan-Meier (KM) curve, the fitted exponential model, and the numbers at risk. A single Bonferroni rejection of an interval is good reasoning for rejecting a model. In Figure 1, the very poor fit (Bonferroni rejection) is at the end of the curve. This is very unfavourable in health economic evaluation, where plausible extrapolations beyond the observed data are paramount to model selection. That is, the Bonferroni rejection at the end of the curve indicates that the extrapolation beyond the trial period may be poor.
@@
An overall p-value for the fitted exponential model is also of interest. For the transformed Fisher test statistic, there are $I=42$ intervals, and our test statistic was computed as $2\sum_{j=1}^{I}log(U_{j})=81.15$, giving a p-value of 0.568. For the PAVSI test statistic, there are $I=42$ intervals. From Figure 1, we can read off that $T_{pavsi}=4$ (as there are 3 individual rejections and 1 Bonferroni rejection). This gives us an overall midpoint p-value of 0.107. (These test statistics should be used with caution as our simulation indicated that they are somewhat conservative in practice.)
@@
Figure 1: Example 1: Plot results. Part a: The KM curve for the data in purple, censors are given by dashed vertical black lines, and the fitted exponential curve is the overlaid curve in black. Part b: The binomial interval test for this data. Intervals are defined by the censors, where for each interval, red indicates a Bonferroni rejection, and grey indicates an individual flag. Part c: Numbers at risk for the data